OpenStudy (anonymous):

The sum of the digits of a two digit number is multiplied by 8. The result is 8 more than the 2-digit number. Find the number

OpenStudy (anonymous):

it can be mentally calculated that the 2-digit number =32

OpenStudy (anonymous):

but let us form equations and solve it

OpenStudy (anonymous):

let the digit in the unit place be x and that in the tenth place be y then we can say the reqd 2-digit number is 10y+x now as per the problem y+x=8 (eq 1) and 8*(x+y) = 10y+x +8 8x+8y=10y+x+8 8x-x +8y -10y =8 7x -2y =8 (eq 2) just solve (eq 1) & (eq 2) get the values of x and y and plug them in 10y+x

OpenStudy (anonymous):

x= 8/3 y= 16/3 The required two-digit number = 10y+x=56

OpenStudy (anonymous):

how do u got 32 answer. @matricked